# Fairness-Aware Domain Generalization under Covariate and Dependence Shifts

### v1

@@ Abstract
Achieving the generalization of an invariant classifier from source domains to shifted target domains while simultaneously considering model fairness is a substantial and complex challenge in machine learning. Existing domain generalization research typically attributes domain shifts to concept shift, which relates to alterations in class labels, and covariate shift, which pertains to variations in data styles. In this paper, by introducing another form of distribution shift, known as dependence shift, which involves variations in fair dependence patterns across domains, we propose a novel domain generalization approach that addresses domain shifts by considering both covariate and dependence shifts. We assert the existence of an underlying transformation model can transform data from one domain to another. By generating data in synthetic domains through the model, a fairness-aware invariant classifier is learned that enforces both model accuracy and fairness in unseen domains. Extensive empirical studies on four benchmark datasets demonstrate that our approach surpasses state-of-the-art methods.
@@ 1 Introduction
While modern fairness-aware machine learning techniques have demonstrated significant success in various applications [1, 2, 3], their primary objective is to facilitate equitable decision-making, ensuring fairness across all demographic groups, regardless of sensitive attributes, such as race and gender. Nevertheless, state-of-the-art methods can encounter severe shortcomings during the inference phase, mainly due to poor generalization when the spurious correlation deviates from the patterns seen in the training data. This correlation can manifest either between model outcomes and sensitive attributes [4, 5] or between model outcomes and non-semantic data features [6]. This issue originates from the existence of out-of-distribution (OOD) data, resulting in catastrophic failures.
Over the past decade, the machine learning community has made significant strides in studying the OOD generalization (or domain generalization, DG) problem and attributing the cause of the poor generalization to the distribution shifts from source domains to target domains. There are two dominant shift types [7]: concept shift and covariate shift. Concept shift refers to OOD samples drawn from a distribution with semantic change e.g., dog v.s. cat, and covariate shift characterizes the extent to which the distributions of data features differ across domains e.g., photo v.s. cartoon. While a variety of DG approaches have been explored, these methods often exhibit two specific limitations: (1) they predominantly address a single type of distribution shift (either concept or covariate shift) and disregard the significance of integrating fairness considerations into the learning process [8, 9, 10, 11, 12, 13], or (2) existing methods assume that the mentioned distribution shifts remain static, and there exist distinct fairness dependence between model outcomes and sensitive attributes in the source and target domains [4, 5]. Recently, Pham et al., [6] address the DG problem with a focus on model fairness and accuracy in the presence of covariate shift. However, this work assumes that the fair dependence patterns across domains remain constant. Hence, there is a need for research that delves into the fairness-aware DG problem in the context of covariate shift and accounts for the varying fairness dependence between source and target domains. Details and more related works are summarized in Appendix A.
In this paper, in addition to concept shift and covariate shift, inspired by [4], we first introduce a novel type of shift known as “dependence shift" for DG problems. The dependence shift is characterized by a disparity in the dependence between sensitive attributes and sample classes within source domains as compared to target domains. As illustrated in Fig. 1, in the source domain 1, images labeled with “Dog" exhibit a strong correlation with the sensitive attribute “Grass", whereas “Cat" images are correlated with “Couch". However, in the target domain, this correlation between image labels and sensitive attributes is reversed and shifted from source domains. Furthermore, we define the fairness-aware DG problem within a broader context, where we account for two types of shifts, covariate shift and dependence shift, that occur when transferring from observed source domains to an unknown and inaccessible target domain.
Figure 1: Illustration of fairness-aware domain generalization problems. Different data domains correspond to different image styles (“Photo", “Cartoon", and “Painting"). Each domain is associated with various fair dependencies between class labels (“Dog" and “Cat") and sensitive attributes (“Grass" and “Couch"), estimated using the demographic parity in Sec. C.3. In the Photo domain, (mostly) dogs in the grass and cats in couches. In the Painting domain, (mostly) dogs in couches and cats in the grass.
@@
As shown in Fig. 1, various image styles (“Photo", “Cartoon" and “Painting") specify different data domains, and “Grass" and “Couch" correspond to sensitive attributes. The goal of the problem is to find a classifier that remains invariant in classifying between “Dog" and “Cat" across the observed source domains and subsequently enhance its generalization performance when faced with a target domain that is unseen during training characterized by different stylistic variations and fairness dependencies. To tackle the problem, we introduce a new framework, which we called Fair Disentangled Domain Generalization (FDDG). The key idea in our framework revolves around understanding transformations that account for both covariate and dependence shifts, enabling the mapping of data between domains, and then subsequently enforcing invariance by generating synthetic domains through these transformations. We leverage this framework to systematically define the fairness-aware DG problem as a semi-infinite constrained optimization problem. We then apply this re-formulation to demonstrate that a tight approximation of the problem can be achieved by solving the empirical, parameterized dual for this problem. Moreover, we develop a novel interpretable bound focusing on fairness within a target domain, considering the DG arising from both covariate and dependence shifts. Finally, extensive experimental results on the proposed new algorithm show that our algorithm significantly outperforms state-of-the-art baselines on several benchmarks.
Contributions. Our main contributions are summarized
@@
  • To our knowledge, we are the first to introduce a fairness-aware DG problem within a framework that accommodates inter-domain variations arising from two distinct types of distribution shifts, covariate shift and dependence shift.
@@
  • We present a novel algorithm designed to address the fairness-aware DG problem. This algorithm enforces invariance across unseen target domains by utilizing generative models derived from the observed source domains.
  • Comprehensive experiments are conducted to verify the effectiveness of FDDG. We empirically show that our algorithm significantly outperforms state-of-the-art baselines on four benchmarks.
@@ 2 Preliminaries
Notations. Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ denote a feature space, $\mathcal{Z}=\{-1,1\}$ is a sensitive label space, and $\mathcal{Y}=\{0,1\}$ is a label space for classification. Let $\mathcal{C}\subseteq\mathbb{R}^{c}$, $\mathcal{A}\subseteq\mathbb{R}^{a}$, and $\mathcal{S}\subseteq\mathbb{R}^{s}$ be the latent content, sensitive and style spaces, respectively, induced from $\mathcal{X}$ and $\mathcal{A}$ by an underlying transformation model $T:\mathcal{X}\times\mathcal{Z}\times\mathcal{E}\rightarrow\mathcal{X}\times\mathcal{Z}$. We use $X,Z,Y,C,A,S$ to denote random variables that take values in $\mathcal{X},\mathcal{Z},\mathcal{Y},\mathcal{C},\mathcal{A},\mathcal{S}$ and $\mathbf{x},z,y,\mathbf{c},\mathbf{a},\mathbf{s}$ the realizations. A domain $e\in\mathcal{E}$ is specified by distribution $\mathbb{P}(X^{e},Z^{e},Y^{e}):\mathcal{X}\times\mathcal{Z}\times\mathcal{Y}\rightarrow[0,1]$. A classifier $f$ in a class space $\mathcal{F}$ denotes $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$.
@@
Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},z_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$ sampled i.i.d. from $\mathbb{P}(X,Z,Y)$, a classifier $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ is fair when the prediction $\hat{Y}=f(X)$ is independent of sensitive random variable $Z$. To get rid of the indicator function and relax the exact values, a linear approximated form of the difference between sensitive subgroups is defined as

### v2

# Waco, Texas

@@ Abstract
The endeavor to preserve the generalization of a fair and invariant classifier across domains, especially in the presence of distribution shifts, becomes a significant and intricate challenge in machine learning. In response to this challenge, numerous effective algorithms have been developed with a focus on addressing the problem of fairness-aware domain generalization. These algorithms are designed to navigate various types of distribution shifts, with a particular emphasis on covariate and dependence shifts. In this context, covariate shift pertains to changes in the marginal distribution of input features, while dependence shift involves alterations in the joint distribution of the label variable and sensitive attributes. In this paper, we introduce a simple but effective approach that aims to learn a fair and invariant classifier by simultaneously addressing both covariate and dependence shifts across domains. We assert the existence of an underlying transformation model can transform data from one domain to another, while preserving the semantics related to non-sensitive attributes and classes. By augmenting various synthetic data domains through the model, we learn a fair and invariant classifier in source domains. This classifier can then be generalized to unknown target domains, maintaining both model prediction and fairness concerns. Extensive empirical studies on four benchmark datasets demonstrate that our approach surpasses state-of-the-art methods. Code repository is available at https://github.com/jk-kaijiang/FDDG.
Keywords fairness  $\cdot$ generalization  $\cdot$ distribution shifts
@@ 1 Introduction
Figure 1: Illustration of the problem in generalizing fair classifiers across different data domains under covariate and dependence shifts simultaneously. (Left) Images in source and target domains have different styles (Photos and Arts). Each data domain is linked to a distinct correlation between class labels (NC and C) and sensitive attributes (Male and Female). (Right) We consider $\mathbf{x}=[x_{1},x_{2}]^{T}$ a simple example of a two-dimensional feature vector. A fair classifier $f$ learned using source data is applied to data sampled from various types of shifted target domains, resulting in misclassification and unfairness. $f^{*}$ represents the true classifier in the target domain.
@@
While modern fairness-aware machine learning techniques have demonstrated significant success in various applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], their primary objective is to facilitate equitable decision-making, ensuring algorithmic fairness across all demographic groups characterized by sensitive attributes, such as race and gender. Nevertheless, the generalization of a fair classifier learned in the source domain to a target domain during inference often demonstrates severe limitations in many state-of-the-art methods. The poor generalization can be attributed to the data distribution shifts from source to target domains, resulting in catastrophic failures.
There are two main lines of data distribution shifts [11]: general and fairness-specific shifts. The former focuses on shifts involving input features and labels. Specifically, covariate shift [12] and label shift [13] refer to variations due to different marginal distributions over feature and class variables, respectively. Concept shift [14] indicates "functional relation change" due to the change amongst the instance-conditional distributions [15]. Moreover, fairness-specific shifts consider additional sensitive attributes and hence place a greater emphasis on ensuring algorithmic fairness. Demographic shift [^1] Dependence shift is named as correlation shift in [16]. [16] refers to certain sensitive population subgroups becoming more or less probable during inference. Dependence shift [11] captures the correlation change between the class variable and sensitive attributes. Within these distribution shifts, a trained fair classifier from source domains is directly influenced and may degrade when adapted to target domains.
To simplify, we narrow the scope of distribution shifts to two prominent ones: covariate shift, which has been extensively investigated in the context of out-of-distribution (OOD) generalization [15, 17], and dependence shift, a topic that has gained attention in recent research. In the illustrative example shown in Fig. 1, the source and target domains exhibit variations stemming from different image styles (Photos and Arts) and correlations between labels (No-cooking and Cooking) and sensitive attributes (Male and Female). Specifically, in the source domain, most males in the kitchen are not cooking, whereas in the target domain, a distinct correlation is observed with most males engaging in cooking. To learn a classifier that is both fair and accurate under such hybrid shifts, a variety of domain generalization approaches have been explored. Predominantly, these methods often exhibit two specific limitations: they (1) address either covariate shift [15, 18, 17] or dependence shift [19, 20], or (2) solely focus on covariate shift but not explicitly indicate the existence of dependence shift [21]. Therefore, there is a need for research that explores the problem of fairness-aware domain generalization (FDG), considering both covariate and dependence shifts simultaneously across source and target domains.
@@
In this paper, we introduce a novel framework, namely Fair disEntangled DOmain geneRAlization (FEDORA). The key idea in our framework revolves around learning a fair and accurate classifier that can generalize from given source domains to target domains, which remain unknown and inaccessible during training. The variations in these domains result from the concurrent presence of covariate and dependence shifts. Notice that, unlike the settings in some works involving covariate shift [22, 23, 24], we assert each domain possesses a distinct data style (Photos and Arts), resulting in an alternation in feature spaces. Technically, we assert the existence of a transformation model that can disentangle input data to a semantic factor that remains invariant across domains, a style factor that characterizes covariate-related information, and a sensitive factor that captures attributes of a sensitive nature. To enhance the generalization of the training classifier and adapt it to unknown target domains, we augment the data by generating them through the transformation model. It utilizes semantic factors associated with various style and sensitive factors sampled from their respective prior distributions. Furthermore, we leverage this framework to systematically define the FDG problem as a semi-infinite constrained optimization problem. Theoretically, we apply this re-formulation to demonstrate that a tight approximation of the problem can be achieved by solving the empirical, parameterized dual for this problem. Moreover, we develop a novel interpretable bound focusing on fairness within a target domain, considering the domain generalization arising from both covariate and dependence shifts. Finally, extensive experimental results on the proposed new algorithm show that our algorithm significantly outperforms state-of-the-art baselines on several benchmarks. Our main contributions are summarized.
@@
  • We introduce a fairness-aware domain generalization problem within a framework that accommodates inter-domain variations arising from covariate and dependence shifts simultaneously. We also give a brief survey by comparing the setting of related works.
@@
  • We present a novel algorithm, FEDORA, that enforces invariance across unseen target domains by utilizing generative models derived from the observed source domains.
  • Comprehensive experiments are conducted to verify the effectiveness of FEDORA. We empirically show that it significantly outperforms state-of-the-art baselines on four benchmarks.
2 Related Works
Table 1: Different Types of Distribution Shifts.
Type of Shifts Notations, $\forall s\in\mathcal{E}_{s}$
Covariate Shift (Cov.) [12] $\mathbb{P}_{X}^{s}\neq\mathbb{P}_{X}^{t}$
Label Shift (Lab.) [13] $\mathbb{P}_{Y}^{s}\neq\mathbb{P}_{Y}^{t}$
Concept Shift (Con.) [14] $\mathbb{P}_{Y|X}^{s}\neq\mathbb{P}_{Y|X}^{t}$
Demographic Shift (Dem.) [16] $\mathbb{P}_{Z}^{s}\neq\mathbb{P}_{Z}^{t}$
Dependence Shift (Dep.) [11] $\mathbb{P}_{Y|Z}^{s}\neq\mathbb{P}_{Y|Z}^{t}$ and $\mathbb{P}_{Z}^{s}=\mathbb{P}_{Z}^{t}$; or, $\mathbb{P}_{Z|Y}^{s}\neq\mathbb{P}_{Z|Y}^{t}$ and $\mathbb{P}_{Y}^{s}=\mathbb{P}_{Y}^{t}$
Hybrid Shift Any combination of the shifts above.
Table 2: An overview of different settings of existing approaches in mitigating unfairness under distribution shifts.
Refs. Distribution Shifts Spaces Change*,   $\forall s\in\mathcal{E}^{s}$ $|\mathcal{E}^{s}|$ Access to Target
Cov. Lab. Con. Dem. Dep. $\mathcal{X}^{s}\neq\mathcal{X}^{t}$ $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$ $\mathcal{Z}^{s}\neq\mathcal{Z}^{t}$
[22, 23, 24] $\bullet$ 1 No
[25] $\bullet$ M No
[26, 27] $\bullet$ 1 Yes
[21] $\bullet$ $\bullet$ M No
[28] $\bullet$ 1 Yes
[29, 30] $\bullet$ 1 Yes
[31, 16] $\bullet$ $\bullet$ 1 Yes
[20] $\bullet$ M No
[19] $\bullet$ 1 No
[11] $\bullet$ 1 Yes
[32] $\bullet$ $\bullet$ 1 Yes
[33] $\bullet$ $\bullet$ 1 Yes
[34] $\bullet$ $\bullet$ $\bullet$ 1 Yes
[35] $\bullet$ $\bullet$ $\bullet$ 1 No
[36] $\bullet$ $\bullet$ $\bullet$ 1 Yes
FEDORA $\bullet$ $\bullet$ $\bullet$ M No
    $\mathcal{Y}^{s}\neq\mathcal{Y}^{t}$ and $\mathcal{Z}^{s}\neq\mathcal{Z}^{t}$ indicate the introduction of new labels and new sensitive attributes. A change in $\mathcal{X}$ denotes a shift in feature variation, such as transitioning from photo images to arts.
Domain generalization. Addressing the challenge of domain shift and the absence of OOD data has led to the introduction of several state-of-the-art methods in the domain generalization field [37, 38, 17, 15]. These methods are designed to enable deep learning models to possess intrinsic generalizability, allowing them to adapt effectively from one or multiple source domains to target domains characterized by unknown distributions [39]. They encompass various techniques, such as aligning source domain distributions to facilitate domain-invariant representation learning [40], subjecting the model to domain shift during training through meta-learning [41], and augmenting data with domain analysis, among others [42], and so on. In the context of the number of source domains, a significant portion of research [17, 15, 43] has focused on the multi-source setting. This setting assumes the availability of multiple distinct but relevant domains for the generalization task. As mentioned in [43], the primary motivation for studying domain generalization is to harness data from multiple sources in order to unveil stable patterns. This entails learning representations invariant to the marginal distributions of data features, all while lacking access to the target data. Nevertheless, existing domain generalization methods tend to overlook the aspect of learning with fairness, where group fairness dependence patterns may not change domains.
Fairness learning for changing environments. Two primary research directions aim to tackle fairness-aware machine learning in dynamic or changing environments. The first approach involves equality-aware monitoring methods [44, 45, 21, 26, 33, 16, 36], which strive to identify and mitigate unfairness in a model’s behavior by continuously monitoring its predictions. These methods adapt the model’s parameters or structure when unfairness is detected. However, a significant limitation of such approaches is their assumption of invariant fairness levels across domains, which may not hold in real-world applications. The second approach [19, 20] focuses on assessing a model’s fairness in a dynamic environment exclusively under dependence shifts. However, it does not consider other types of distribution shifts.
In response to these limitations, this paper adopts a novel approach by attributing the distribution shift from source to target domains to both covariate shift and fairness dependence shift simultaneously. The objective is to train a fairness-aware invariant classifier capable of effective generalization across domains, ensuring robust performance in terms of both model accuracy and the preservation of fair dependence between predicted outcomes and sensitive attributes under both shifts.
@@ 3 Preliminaries
Notations. Let $\mathcal{X}\subseteq\mathbb{R}^{d}$ denote a feature space, $\mathcal{Z}=\{-1,1\}$ is a sensitive space, and $\mathcal{Y}=\{0,1\}$ is a label space for classification. Let $\mathcal{C}\subseteq\mathbb{R}^{c}$, $\mathcal{A}\subseteq\mathbb{R}^{a}$, and $\mathcal{S}\subseteq\mathbb{R}^{s}$ be the semantic, sensitive and style latent spaces, respectively, induced from $\mathcal{X}$ and $\mathcal{A}$ by an underlying transformation model $T:\mathcal{X}\times\mathcal{Z}\times\mathcal{E}\rightarrow\mathcal{X}\times\mathcal{Z}$. We use $X,Z,Y,C,A,S$ to denote random variables that take values in $\mathcal{X},\mathcal{Z},\mathcal{Y},\mathcal{C},\mathcal{A},\mathcal{S}$ and $\mathbf{x},z,y,\mathbf{c},\mathbf{a},\mathbf{s}$ the realizations. A domain $e\in\mathcal{E}$ is defined as a joint distribution $\mathbb{P}^{e}_{XZY}=\mathbb{P}(X^{e},Z^{e},Y^{e}):\mathcal{X}\times\mathcal{Z}\times\mathcal{Y}\rightarrow[0,1]$. A classifier $f$ in a class space $\mathcal{F}$ denotes $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$. We denote $\mathcal{E}$ and $\mathcal{E}^{s}\subset\mathcal{E}$ as the set of domain labels for all domains and source domains, respectively. Superscripts in the samples denote their domain labels, while subscripts specify the indices of encoders. For example, $E_{s}(\mathbf{x}^{s})$ denotes a sample $\mathbf{x}$ drawn from the $s$ domain and encoded by a style encoder $E_{s}$.
@@
Given a dataset $\mathcal{D}=\{(\mathbf{x}_{i},z_{i},y_{i})\}_{i=1}^{|\mathcal{D}|}$ sampled i.i.d. from $\mathbb{P}_{XZY}$, a classifier $f\in\mathcal{F}:\mathcal{X}\rightarrow\mathcal{Y}$ is fair when the prediction $\hat{Y}=f(X)$ is independent of the sensitive random variable $Z$. To get rid of the indicator function and relax the exact values, a linear approximated form of the difference between sensitive subgroups is defined as
$\rho(\hat{Y},Z)=\big{|}\mathbb{E}_{\mathbb{P}_{XZY}}g(\hat{Y},Z)\big{|}\quad\text{where}\quad g(\hat{Y},Z)=\frac{1}{p_{1}(1-p_{1})}\Big{(}\frac{Z+1}{2}-p_{1}\Big{)}\hat{Y}$ (1)
Specifically, when $p_{1}=\mathbb{P}(Z=1)$ and $p_{1}=\mathbb{P}(Z=1,Y=1)$, the fairness notion $\rho(\hat{Y},Z)$ is defined as the difference of demographic parity and the difference of equalized opportunity, respectively [47]. In this paper, we will present the results under demographic parity (and then the expectation in Eq. 1 is over $XZ$), while the framework can be generalized to multi-class, multi-sensitive attributes and other fairness notions. Strictly speaking, a classifier $f$ is fair over subgroups if it satisfies $\rho(\hat{Y},Z)=0$.
Figure 2: (Left) A transformation model $T$ is trained using a bi-directional reconstruction loss (data reconstruction and factor reconstruction) and a sensitiveness loss. (Right) To enhance the generalization of the classifier $f$ to unseen target domains, the transformation model $T$ is used for augmentation in synthetic domains by generating data based on invariant semantic factors and randomly sampled sensitive and style factors that encode synthetic domains. We demonstrate the concept using the ccMNIST dataset, where the domains are distinguished by different digit colors and fair dependencies between class labels and sensitive attributes. Here, sensitive attributes are defined by image background colors.
Problem setting. Given a dataset $\mathcal{D}=\{\mathcal{D}^{e}\}_{e=1}^{|\mathcal{E}|}$, where each $\mathcal{D}^{e}=\{(\mathbf{x}^{e}_{i},z^{e}_{i},y^{e}_{i})\}_{i=1}^{|\mathcal{D}^{e}|}$ is i.i.d. sampled from a domain $\mathbb{P}^{e}_{XZY}$ and $e\in\mathcal{E}$, we consider multiple source domains $\{\mathbb{P}^{s}_{XZY}\}_{s=1}^{|\mathcal{E}^{s}|}$ and a distinct target domain $\mathbb{P}^{t}_{XZY},t\neq s,\forall s\in\mathcal{E}^{s}\subset\mathcal{E}$ and $t\in\mathcal{E}\backslash\mathcal{E}^{s}$, which is unknown and inaccessible during training. Given samples $\{\mathcal{D}^{s}\}_{s=1}^{|\mathcal{E}^{s}|}$ from finite source domains, the goal of fairness-aware domain generalization problems is to learn a classifier $f\in\mathcal{F}$ that is generalizable across all possible domains.
Problem 1 (Fairness-aware Domain Generalization).
Let $\{\mathbb{P}^{s}_{XZY}\}_{s=1}^{|\mathcal{E}^{s}|}$ be a finite subset of source domains and assume that, for each $s\in\mathcal{E}^{s}$, we have access to its corresponding dataset $\mathcal{D}^{s}=\{(\mathbf{x}^{s}_{i},z^{s}_{i},y^{s}_{i})\}_{i=1}^{|\mathcal{D}^{s}|}$ sampled i.i.d from $\mathbb{P}^{s}_{XZY}$. Given a classifier set $\mathcal{F}$ and a loss function $\ell:\mathcal{Y\times Y}\rightarrow\mathbb{R}$, the goal is to learn a fair classifier $f\in\mathcal{F}$ for any $\mathcal{D}^{s}$ that minimizes the worst-case risk over all domains in $\{\mathbb{P}^{e}_{XZY}\}_{e=1}^{|\mathcal{E}|}$ satisfying a group fairness constraint:
$\min_{f\in\mathcal{F}}\>\max_{e\in\mathcal{E}}\>\mathbb{E}_{\mathbb{P}^{s}_{XZY}}\ell(f(X^{s}),Y^{s}),\>\>\>\text{s.t.}\>\>\rho(f(X^{s}),Z^{s})=0$ (2)
The goal of Prob. 1 is to seek a fair classifier $f$ that generalizes from the given finite set of source domains to give a good generalization performance on all domains. Since we do not assume data from a target domain is accessible, it makes Prob. 1 challenging to solve.
Another challenge is how closely the data distributions in unknown target domains match those in the observed source domains. As discussed in Sec. 1 and Tab. 1, there are five different types of distribution shifts. In this paper, we narrow the scope and claim the shift between source and target domains is solely due to covariate and dependence shifts.
Definition 2 (Covariate Shift [15] and Dependence Shift[11]).
In Prob. 1, covariate shift occurs when domain variation is attributed to disparities in the marginal distributions over input features $\mathbb{P}^{s}_{X}\neq\mathbb{P}^{t}_{X},\forall s$. On the other hand, Prob. 1 exhibits a dependence shift when domain variation arises from alterations in the joint distribution between $Y$ and $Z$, denoted $\mathbb{P}^{s}_{YZ}\neq\mathbb{P}^{t}_{YZ},\forall s$ where $\mathbb{P}_{Y|Z}^{s}\neq\mathbb{P}_{Y|Z}^{t}$ and $\mathbb{P}_{Z}^{s}=\mathbb{P}_{Z}^{t}$; or $\mathbb{P}_{Z|Y}^{s}\neq\mathbb{P}_{Z|Y}^{t}$ and $\mathbb{P}_{Y}^{s}=\mathbb{P}_{Y}^{t}$.
Underlying transformation models. Inspired by existing domain generalization endeavors [15, 17, 48], distribution shifts can characterize generalization tasks across domains through an underlying transformation model $T$. The motivation behind using $T$ lies in bolstering the robustness and adaptability of the classifier $f$ across diverse domains. By learning a transformation model, the objective is twofold: (1) to enable the model to adapt domain-invariant data representations (factors) from the input data by disentangling domain-specific variations and (2) to generate augmented data in new domains by perturbing existing samples with various variations. This augmentation enhances the diversity of the source data and thereby improves the ability to generalize to unseen target domains.
4 Methodology
4.1 Learning the Transformation Model
One goal of the transformation model $T=\{E,G\}$ is to disentangle an input sample from source domains into three factors in latent spaces by learning a set of encoder $E=\{E_{c},E_{a},E_{s}\}$ and a decoder $G:\mathcal{C}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathcal{X}$, where $E_{c}:\mathcal{X}\rightarrow\mathcal{C}$, $E_{a}:\mathcal{X}\rightarrow\mathcal{A}$, and $E_{s}:\mathcal{X}\rightarrow\mathcal{S}$ represent semantic, sensitive and style encoders, respectively.
Assumption 1 (Multiple Latent Factors).
Given dataset $\mathcal{D}^{e}=\{(\mathbf{x}^{e}_{i},z^{e}_{i},y^{e}_{i})\}_{i=1}^{|\mathcal{D}^{e}|}$ sampled i.i.d. from $\mathbb{P}^{e}_{XZY}$ domain $e\in\mathcal{E}$, we assume that each instance $\mathbf{x}^{e}_{i}$ is generated from (1) a latent semantic factor $\mathbf{c}\in\mathcal{C}$, where $\mathcal{C}=\{\mathbf{c}_{y=0},\mathbf{c}_{y=1}\}$; (2) a latent sensitive factor $\mathbf{a}\in\mathcal{A}$, where $\mathcal{A}=\{\mathbf{a}_{z=1},\mathbf{a}_{z=-1}\}$; and (3) a latent style factor $\mathbf{s}^{e}$, where $\mathbf{s}^{e}$ is specific to the individual domain $e$. We assume that the semantic and sensitive factors in $\mathcal{C}$ and $\mathcal{A}$ do not change across domains. Each domain $\mathbb{P}^{e}_{XZY}$ is represented by a style factor $\mathbf{s}^{e}$ and the dependence score $\rho^{e}=\rho(Y^{e},Z^{e})$ [^2] Here, $\rho$ functions equivalently as it does in Eq. 1, by substituting $\hat{Y}$ to $Y$., denoted $e:=(\mathbf{s}^{e},\rho^{e})$, where $\mathbf{s}^{e}$ and $\rho^{e}$ are unique to the domain $\mathbb{P}^{e}_{XZY}$.
Note that Assump. 1 is similarly related to the one made in [17, 15, 48, 49]. In our paper, with a focus on group fairness, we expand upon the assumptions of existing works by introducing three latent factors. Under Assump. 1, if two instances $(\mathbf{x}^{e_{i}},z^{e_{i}},y)$ and $(\mathbf{x}^{e_{j}},z^{e_{j}},y)$ where $e_{i},e_{j}\in\mathcal{E},i\neq j$ share the same class label, then the latter instance can be reconstructed by decoder $G$ from the former using $\mathbf{c}=E_{c}(\mathbf{x}^{e_{i}})$, $\mathbf{s}=E_{s}(\mathbf{x}^{e_{j}})$, and $\mathbf{a}=E_{a}(\mathbf{x}^{e_{j}})$ through $T$, denoted $(\mathbf{x}^{e_{j}},z^{e_{j}})=T(\mathbf{x}^{e_{i}},z^{e_{i}},e_{j})$.
To enhance the effectiveness of the transformation model $T$, our overall learning loss for these encoders and decoders consists of two main components: a bidirectional reconstruction loss and a sensitiveness loss.